# The basis problem for modular forms for the Weil representation

### v1

@@ 4 Vector valued Eisenstein series
5 Vector valued Siegel–Weil formula
6 Space of theta series
@@ 1 Introduction
Let $L$ be a positive-definite even lattice of even rank $m$ with bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). It is well-known that the theta series $\theta(\tau)=\sum_{\lambda\in L}e^{\pi i(\lambda,\lambda)\tau}$ of $L$ is a modular form of weight $m/2$ for the congruence subgroup $\Gamma_{0}(N)$ and a certain quadratic character, where $N$ is the level of $L$. In [E1] Eichler announced and later in [E2] proved that for $N$ square-free and $m=0\bmod 4$, the space of new forms of weight $m/2$ for $\Gamma_{0}(N)$ has a basis consisting of theta series corresponding to lattices of level $N$. More generally the question of finding an explicit basis of an appropriate space of modular forms consisting of theta series is known as the basis problem. An important contribution to this problem was given by Waldspurger. In [Wal] he proved that for all positive integers $m=0\bmod 4$ and $k\geq m/2$ the space of new forms of weight $k$ for $\Gamma_{0}(N)$ and trivial character is generated by theta series of lattices of rank $m$ and level $N$ weighted with harmonic polynomials of degree $k-m/2$. For non-trivial character he proved the result when $N$ is a square free integer congruent to $1$ modulo $4$. Furthermore, in [BKS] Böcherer, Katsurada and Schulze-Pillot showed that for $k>2n+1$ the space of Siegel cusp forms of weight $k$ and genus $n$ for $\Gamma_{0}(N)$ with $N$ square free is generated by harmonic theta series of appropriate lattices.
@@
$\theta_{L,P}(M\tau)=(c\tau+d)^{k}\rho_{L^{\prime}/L}(M)\theta_{L}(\tau)\quad\text{for $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\mathrm{SL}_{2}(\mathbb{Z})$},$
where $M\tau$ denotes the Möbius transformation and $\rho_{L^{\prime}/L}:\mathrm{SL}_{2}(\mathbb{Z})\rightarrow\operatorname{GL}(\mathbb{C}[L^{\prime}/L])$ the Weil representation of $\mathrm{SL}_{2}(\mathbb{Z})$. The latter is a special case of the representations of symplectic groups constructed by Weil in [W1]. Two positive-definite even lattices $L$ and $M$ of rank $m$ have isomorphic discriminant forms $\mathcal{D}$ and hence isomorphic Weil representations if and only if they are in the same genus, which is denoted by $I\!I_{m,0}(\mathcal{D})$. It is a natural question whether the space $\mathrm{S}_{k}(\mathcal{D})$ of cusp forms for $\rho_{\mathcal{D}}$ is generated by the theta series in the genus $I\!I_{m,0}(\mathcal{D})$. The present paper answers this question in the affirmative if the rank of the lattice is sufficiently large compared to the $p$-ranks of the discriminant form. Because there is no canonical isomorphism between discriminant forms of lattices in the same genus, there is also no canonical way to identify their Weil representations. Therefore, we define for a discriminant form $\mathcal{D}$
@@
in the notation of Conway and Sloane (cf. [CS], chapter 15) and $|\mathcal{D}|=p^{\alpha}a$ with $(a,p)=1$.
@@
Corollary 6.8 follows immediately from Theorem 6.7 using a result in [] (see end of Section 5).
The proof of Theorem 6.7 uses the so-called doubling method that was also employed in [BKS]. In what follows we give a short description of the proof idea.
@@
I would like to thank my advisor N. Scheithauer for suggesting this topic as part of my doctoral thesis and for the support he has given me in pursuing it. Moreover, I would like to thank J. H. Bruinier, P. Kiefer, I. Metzler and R. Zufetti for stimulating discussions and helpful comments.
@@ 2 Weil representation
References for discriminant forms are [AGM], [Bo3], [CS], [N], [S1] and [Sk].
@@
is known as Milgram’s formula. See [S1] for more general Gauss sums.
@@
Now we want to recall some facts about the symplectic group. A nice reference is [F].
@@
Let $\mathcal{D}$ be a discriminant form of even signature and level $N$ and let $\mathbb{C}[\mathcal{D}^{n}]$ be the group algebra of $\mathcal{D}^{n}=\mathcal{D}\times\ldots\times\mathcal{D}$ spanned by a formal basis $(e^{\underline{\gamma}})_{\underline{\gamma}\in\mathcal{D}^{n}}$, where $\underline{\gamma}=(\gamma_{1},\ldots,\gamma_{n})$. Then the Weil representation $\rho_{\mathcal{D}}^{(n)}:\mathrm{Sp}_{2n}(\mathbb{Z})\to\mathbb{C}[\mathcal{D}^{n}]$ of $\mathrm{Sp}_{2n}(\mathbb{Z})$ can be defined by (cf. [W1], [Bo1] and [Zh])
@@
Then the Weil representation is unitary with respect to this scalar product. There is a natural isomorphism $\mathbb{C}[\mathcal{D}^{n}]\cong\mathbb{C}[\mathcal{D}]^{\otimes n}$ that we will make frequent use of. For the following cf. [St, Lemma 3.4].
@@
(cf. for example [Br], [S2] or [S3]). The following result is easy to prove.
@@
We will now define modular forms for the Weil representation of $\mathrm{Sp}_{2n}(\mathbb{Z})$ generalizing the definition from [F] to the vector valued case (cf. also [Br] and [Zh]). Let
@@
be the Siegel upper half space. Let $k\in\mathbb{Z}$ and $f$ be a function from $\mathbb{H}$ to a complex vector space. Let $M=\left(\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right)\in\operatorname{GL}_{2n}(\mathbb{R})$ be a symplectic similitude matrix, i.e. $M^{T}J_{n}M=lJ_{n}$ for some $l\in\mathbb{R}$. We define the Petersson-slash operator $|_{k}$ by
@@
We now want to define theta series weighted with harmonic polynomials. Again we generalize the definition in [F] to the vector valued case in the same way as was done for $n=1$ for example in [Bo1].
@@
A harmonic form of degree $h$ in the matrix variable $X=(x_{ij})\in\mathbb{C}^{m,n}$ is a complex polynomial $P(X)$ with the properties
@@
The genus $G$ of an even lattice $L$ is the set of isometry classes of lattices equivalent to $L$ over $\mathbb{Z}_{p}$ for all primes $p$ and over $\mathbb{R}$. It is uniquely determined by the signature $(t_{+},t_{-})$ of $L$ and its discriminant form $L^{\prime}/L$ (cf. [N]). Therefore, the positive-definite even lattices $L$ of rank $m$ with $L^{\prime}/L\cong\mathcal{D}$ form a genus denoted by $I\!I_{m,0}(\mathcal{D})$.
@@
$\vartheta_{G,k}\coloneqq\partial_{h}\theta_{G}^{(2)}\in\Theta_{m,m/2+h}(\mathcal{D})\otimes\Theta_{m,m/2+h}(\mathcal{D}).$
Note that $\partial_{h}$ is essentially the operator $\mathcal{D}_{h}$ defined by Eichler and Zagier in [EZ]. It is a special case of the operators studied in [I].
@@
An element $S\in\operatorname{SO}(m)$ naturally acts on $\operatorname{H}_{m}^{h}$ by $S.p=p(S^{T}\cdot)$ and it is well-known that this representation is irreducible (see for example [KV, (0.9) and (5.7)]). Note that we have $\nabla S.p=S\cdot(\nabla p)(S^{T}\cdot)$ so that for any $p,q\in\operatorname{H}_{m}^{h}$ we obtain
@@ 3 Vector valued Hecke operators
We now want to define vector valued Hecke operators. They were introduced by Bruinier and Stein in [BS] and naturally appear when we use the doubling method. We will study some of their properties and describe their kernel functions. In this section we will only consider the case $n=1$. So from now on, unless stated otherwise, whenever $n$ is omitted, it is assumed to be equal to $1$.
Let $\mathcal{D}$ be a discriminant form of even signature and level $N$. We define Hecke operators $T(l^{2})$ acting on $\mathrm{M}_{k}(\mathcal{D})$ (see [BS]). In order to do so we extend the right action given by the inverse of the Weil representation to certain matrices in $\mathrm{Mat}_{2}(\mathbb{Z})$. Let $l$ be a non-negative integer and $\alpha=\left(\begin{smallmatrix}l^{2}&0\\ 0&1\end{smallmatrix}\right)\in\mathrm{Mat}_{2}(\mathbb{Z})$. We define
@@
It is shown in [BS] that $\rho_{\mathcal{D}}(\delta)^{-1}$ is well-defined and that
@@
Here $\operatorname{oddity}(\mathcal{D})\bmod 8$ depends only on the $2$-adic subgroup of $\mathcal{D}$ and is even if the signature of $\mathcal{D}$ is even. For a precise definition see [CS, Chapter 15]. The following lemma is well-known (see e.g. [F, Hilfssatz IV.1.12])
@@
The operators $T(l^{2})$ with $l$ is coprime to $N$ behave analogously to the classical Hecke operators for $\mathrm{SL}_{2}(\mathbb{Z})$.
@@
There exists a basis of $\mathrm{S}_{k}(\mathcal{D})$ consisting of simultaneous eigenforms for all Hecke operators $T(l^{2})$ with $(l,N)=1$. Let $f$ be a simultaneous eigenform with eigenvalues $\lambda(l^{2})$. The $L$-series
@@
and so this algebra is a commutative algebra of self-adjoint operators. (More details of this Hecke algebra can be found in the upcoming [Me].) This implies that $\mathrm{S}_{k}(\mathcal{D})$ has a basis consisting of simultaneous eigenforms. Let $f$ be an eigenform with eigenvalues $\lambda(l^{2})$ for $T(l^{2})$ and $\widetilde{\lambda}(l^{2})$ for $\widetilde{T}(l^{2})$. We define
@@
where $T^{\Gamma(N)}(l^{2})$ is the standard Hecke operator on $\mathrm{S}_{k}(\Gamma(N))$ (cf. [Wr, Theorem 31 (ii)]). Since the elements of the form $v_{\gamma,\chi}$ generate $\mathbb{C}[\mathcal{D}]$, we find a pair $(\gamma,\chi)$ such that
@@
and $g\in\mathrm{S}_{k}(\Gamma(N))$ is a simultaneous eigenform with eigenvalues $\chi(l)^{-1}\widetilde{\lambda}(l^{2})$. Now the convergence of $\widetilde{L}(f,s)$ follows from the scalar case (see for example [R]). Again we write
@@
Let $\mathcal{D}$ be a discriminant form, $p$ a prime and $r\geq 1$. Let $f\in\mathrm{M}_{k}(\mathcal{D})$ and $\gamma\in\mathcal{D}$ with $\gamma\not\in\mathcal{D}^{p}$ and if $p=2$, also $\gamma\not\in\mathcal{D}^{2*}$. Then
@@
Let $0<s<2r$. Similar to $s=2r$, using the explicit formula for the action of an arbitrary element in the Weil representation from [S1, Theorem 4.7] we obtain
@@
Let $\mathcal{D}$ be a discriminant form of even signature and level $N$, $p$ a prime dividing $N$ and $s\in\mathbb{C}$. Let $f\in\mathrm{S}_{k}(\mathcal{D})$ and assume that
$\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{rs}}\right)f=0.$
Then
$\langle f,e^{\gamma}\rangle=\langle f,e^{\mu}\rangle,$
whenever $p\gamma=p\mu$ and $\operatorname{q}(\gamma)=\operatorname{q}(\mu)\bmod 1$ and $\gamma,\mu\not\in\mathcal{D}^{p}$ and if $p=2$, also $\gamma,\mu\not\in\mathcal{D}^{2*}$.
@@
Since both $\gamma$ and $\mu$ satisfy the condition of Proposition 3.4, we have for $r\geq 1$
$\langle T(p^{2r})f,e^{\gamma}\rangle$ $=T^{\operatorname{q}(\gamma)}(p^{2r})\langle f,e^{p^{r}\gamma}\rangle\text{ and}$
$\langle T(p^{2r})f,e^{\mu}\rangle$ $=T^{\operatorname{q}(\mu)}(p^{2r})\langle f,e^{p^{r}\mu}\rangle.$
Furthermore, $p\gamma=p\mu$ and $\operatorname{q}(\gamma)=\operatorname{q}(\mu)\bmod 1$, so that the right-hand sides are equal, thus
$\langle f,e^{\gamma}\rangle$ $=-\langle\sum_{r=1}^{\infty}\frac{T(p^{2r})}{p^{rs}}f,e^{\gamma}\rangle=-\sum_{r=1}^{\infty}\frac{T^{\operatorname{q}(\gamma)}(p^{2r})}{p^{rs}}\langle f,e^{p^{r}\gamma}\rangle=-\langle\sum_{r=1}^{\infty}\frac{T(p^{2r})}{p^{rs}}f,e^{\mu}\rangle=\langle f,e^{\mu}\rangle.$
@@
The case $x\in\mathbb{Z}$ is well-known (see e.g. [Mi, (7.1.9)]). For $x=\frac{p}{q}$ with $(p,q)=1$ write
@@
for elements $v,w,u\in\mathbb{C}[\mathcal{D}]$ and extend antilinearly in the second argument. The following proposition adapts [Za, Proposition 1], which is originally due to Petersson, to the vector valued case.
@@
where $P_{\gamma,r}$ is the Poincaré series of index $(\gamma,r)$ as defined in [Br]. There it is shown that

### v2

@@ 4 Vector valued Eisenstein series
5 The vector valued Siegel–Weil formula
6 The space of theta series
@@ 1 Introduction
Let $L$ be a positive-definite even lattice of even rank $m$ with bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). It is well-known that the theta series $\theta(\tau)=\sum_{\lambda\in L}e^{\pi i(\lambda,\lambda)\tau}$ of $L$ is a modular form of weight $m/2$ for the congruence subgroup $\Gamma_{0}(N)$ and a certain quadratic character, where $N$ is the level of $L$. In [11] Eichler announced and later in [12] proved that for $N$ square-free and $m=0\bmod 4$, the space of newforms of weight $m/2$ for $\Gamma_{0}(N)$ has a basis consisting of theta series corresponding to lattices of level $N$. More generally the question of finding an explicit basis of an appropriate space of modular forms consisting of theta series is known as the basis problem. An important contribution to this problem was given by Waldspurger. In [33] he proved that for all positive integers $m=0\bmod 4$ and $k\geq m/2$ the space of newforms of weight $k$ for $\Gamma_{0}(N)$ and trivial character is generated by theta series of lattices of rank $m$ and level $N$ weighted with harmonic polynomials of degree $k-m/2$. For non-trivial character he proved the result when $N$ is a square free integer congruent to $1$ modulo $4$. Furthermore, in [3] Böcherer, Katsurada and Schulze-Pillot showed that for $k>2n+1$ the space of Siegel cusp forms of weight $k$ and genus $n$ for $\Gamma_{0}(N)$ with $N$ square free is generated by harmonic theta series of appropriate lattices.
@@
$\theta_{L,P}(M\tau)=(c\tau+d)^{k}\rho_{L^{\prime}/L}(M)\theta_{L,P}(\tau)\quad\text{for $M=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\mathrm{SL}_{2}(\mathbb{Z})$},$
where $M\tau$ denotes the Möbius transformation and $\rho_{L^{\prime}/L}:\mathrm{SL}_{2}(\mathbb{Z})\rightarrow\operatorname{GL}(\mathbb{C}[L^{\prime}/L])$ the Weil representation of $\mathrm{SL}_{2}(\mathbb{Z})$. The latter is a special case of the representations of symplectic groups constructed by Weil in [34]. Two positive-definite even lattices $L$ and $M$ of rank $m$ have isomorphic discriminant forms $\mathcal{D}$ and hence isomorphic Weil representations if and only if they are in the same genus, which is denoted by $I\!I_{m,0}(\mathcal{D})$. It is a natural question whether the space $\mathrm{S}_{k}(\mathcal{D})$ of cusp forms for $\rho_{\mathcal{D}}$ is generated by the theta series in the genus $I\!I_{m,0}(\mathcal{D})$. The present paper answers this question in the affirmative if the rank of the lattice is sufficiently large compared to the $p$-ranks of the discriminant form. Because there is no canonical isomorphism between discriminant forms of lattices in the same genus, there is also no canonical way to identify their Weil representations. Therefore, we define for a discriminant form $\mathcal{D}$
@@
in the notation of Conway and Sloane (cf. [10], chapter 15) and $|\mathcal{D}|=p^{\alpha}a$ with $(a,p)=1$.
@@
Corollary 6.8 follows immediately from Theorem 6.7 using a result in [24] (see end of Section 6).
The proof of Theorem 6.7 uses the so-called doubling method that was also employed in [3]. In what follows we give a short description of the proof idea.
@@
I would like to thank my advisor N. Scheithauer for suggesting this topic as part of my doctoral thesis and for the support he has given me in pursuing it. Moreover, I would like to thank J. H. Bruinier, P. Kiefer, I. Metzler and R. Zufetti for stimulating discussions and helpful comments. I also thank C. Huang for pointing out an error in Lemma 6.5 in an earlier version of this paper, which has been corrected in the present version.
@@ 2 The Weil representation
References for discriminant forms are [1], [6], [10], [25], [27] and [31].
@@
is known as Milgram’s formula. See [27] for more general Gauss sums.
@@
Now we want to recall some facts about the symplectic group. A nice reference is [14].
@@
Let $\mathcal{D}$ be a discriminant form of even signature and level $N$ and let $\mathbb{C}[\mathcal{D}^{n}]$ be the group algebra of $\mathcal{D}^{n}=\mathcal{D}\times\ldots\times\mathcal{D}$ spanned by a formal basis $(e^{\underline{\gamma}})_{\underline{\gamma}\in\mathcal{D}^{n}}$, where $\underline{\gamma}=(\gamma_{1},\ldots,\gamma_{n})$. Then the Weil representation $\rho_{\mathcal{D}}^{(n)}:\mathrm{Sp}_{2n}(\mathbb{Z})\to\mathbb{C}[\mathcal{D}^{n}]$ of $\mathrm{Sp}_{2n}(\mathbb{Z})$ can be defined by (cf. [34], [4] and [38])
@@
Then the Weil representation is unitary with respect to this scalar product. There is a natural isomorphism $\mathbb{C}[\mathcal{D}^{n}]\cong\mathbb{C}[\mathcal{D}]^{\otimes n}$ that we will make frequent use of. For the following cf. [32, Lemma 3.4].
@@
(cf. for example [7], [28] or [29]). The following result is easy to prove.
@@
We will now define modular forms for the Weil representation of $\mathrm{Sp}_{2n}(\mathbb{Z})$ generalizing the definition from [14] to the vector valued case (cf. also [7] and [38]). Let
@@
be the Siegel upper half space. Let $k\in\mathbb{Z}$ and $f$ be a function from $\mathbb{H}$ to a complex vector space. Let $M=\left(\begin{smallmatrix}A&B\\ C&D\end{smallmatrix}\right)\in\operatorname{GL}_{2n}(\mathbb{R})$ be a symplectic similitude matrix, i.e. $M^{T}J_{n}M=lJ_{n}$ for some $l\in\mathbb{R}_{>0}$. We define the Petersson-slash operator $|_{k}$ by
@@
We now want to define theta series weighted with harmonic polynomials. Again we generalize the definition in [14] to the vector valued case in the same way as was done for $n=1$ for example in [4].
@@
A harmonic form of degree $h$ in the matrix variable $X=(x_{ij})_{i=1,\ldots,m;j=1,\ldots,n}$ is a complex polynomial $P(X)$ with the properties
@@
The genus $G$ of an even lattice $L$ is the set of isometry classes of lattices equivalent to $L$ over $\mathbb{Z}_{p}$ for all primes $p$ and over $\mathbb{R}$. It is uniquely determined by the signature $(t_{+},t_{-})$ of $L$ and its discriminant form $L^{\prime}/L$ (cf. [25]). Therefore, the positive-definite even lattices $L$ of rank $m$ with $L^{\prime}/L\cong\mathcal{D}$ form a genus denoted by $I\!I_{m,0}(\mathcal{D})$.
@@
$\vartheta_{G,m/2+h}\coloneqq\partial_{h}\theta_{G}^{(2)}\in\Theta_{m,m/2+h}(\mathcal{D})\otimes\Theta_{m,m/2+h}(\mathcal{D}).$
Note that $\partial_{h}$ is essentially the operator $\mathcal{D}_{h}$ defined by Eichler and Zagier in [13]. It is a special case of the operators studied in [16].
@@
An element $S\in\operatorname{SO}(m)$ naturally acts on $\operatorname{H}_{m}^{h}$ by $S.p=p(S^{T}\cdot)$ and it is well-known that this representation is irreducible (see for example [17, (0.9) and (5.7)]). Note that we have $\nabla S.p=S\cdot(\nabla p)(S^{T}\cdot)$ so that for any $p,q\in\operatorname{H}_{m}^{h}$ we obtain
@@ 3 Vector valued Hecke operators
We now want to define vector valued Hecke operators. They were introduced by Bruinier and Stein in [9] and naturally appear when we use the doubling method. We will study some of their properties and describe their kernel functions. In this section we will only consider the case $n=1$. So from now on, unless stated otherwise, whenever $n$ is omitted, it is assumed to be equal to $1$.
Let $\mathcal{D}$ be a discriminant form of even signature and level $N$. We define Hecke operators $T(l^{2})$ acting on $\mathrm{M}_{k}(\mathcal{D})$ (see [9]). In order to do so we extend the right action given by the inverse of the Weil representation to certain matrices in $\mathrm{Mat}_{2}(\mathbb{Z})$. Let $l$ be a non-negative integer and $\alpha=\left(\begin{smallmatrix}l^{2}&0\\ 0&1\end{smallmatrix}\right)\in\mathrm{Mat}_{2}(\mathbb{Z})$. We define
@@
It is shown in [9] that $\rho_{\mathcal{D}}(\delta)^{-1}$ is well-defined and that
@@
Here $\operatorname{oddity}(\mathcal{D})\bmod 8$ depends only on the $2$-adic subgroup of $\mathcal{D}$ and is even if the signature of $\mathcal{D}$ is even. For a precise definition see [10, Chapter 15]. The following lemma is well-known (see e.g. [14, Hilfssatz IV.1.12])
@@
The operators $T(l^{2})$ with $l$ coprime to $N$ behave analogously to the classical Hecke operators for $\mathrm{SL}_{2}(\mathbb{Z})$.
@@
The algebra generated by all Hecke operators $T(l^{2})$ with $(l,N)=1$ is a commutative algebra of self-adjoint operators. Hence, there exists a basis of $\mathrm{S}_{k}(\mathcal{D})$ consisting of simultaneous eigenforms for it. Let $f$ be a simultaneous eigenform with eigenvalues $\lambda(l^{2})$. The $L$-series
@@
and so this algebra is a commutative algebra of self-adjoint operators. (More details of this Hecke algebra can be found in the upcoming [22].) This implies that $\mathrm{S}_{k}(\mathcal{D})$ has a basis consisting of simultaneous eigenforms. Let $f$ be an eigenform with eigenvalues $\lambda(l^{2})$ for $T(l^{2})$ and $\widetilde{\lambda}(l^{2})$ for $\widetilde{T}(l^{2})$. We define
@@
where $T^{\Gamma(N)}(l^{2})$ is the standard Hecke operator on $\mathrm{S}_{k}(\Gamma(N))$ (cf. [36, Theorem 31 (ii)]). Since the elements of the form $v_{\gamma,\chi}$ generate $\mathbb{C}[\mathcal{D}]$, we find a pair $(\gamma,\chi)$ such that
@@
and $g\in\mathrm{S}_{k}(\Gamma(N))$ is a simultaneous eigenform with eigenvalues $\chi(l)^{-1}\widetilde{\lambda}(l^{2})$. Now the convergence of $\widetilde{L}(f,s)$ follows from the scalar case (see for example [26]). Again we write
@@
Let $\mathcal{D}$ be a discriminant form of even signature, $p$ a prime and $r\geq 1$. Let $f\in\mathrm{M}_{k}(\mathcal{D})$ and $\gamma\in\mathcal{D}$ with $\gamma\not\in\mathcal{D}^{p}$ and if $p=2$, also $\gamma\not\in\mathcal{D}^{2*}$. Then
@@
Let $0<s<2r$. Similar to $s=2r$, using the explicit formula for the action of an arbitrary element in the Weil representation from [27, Theorem 4.7] we obtain
@@
Let $\gamma,\mu\in\mathcal{D}$ and $p$ be a prime. We call the projection of $\gamma$ to the $p$-adic component of $\mathcal{D}$ the $p$-adic component of $\gamma$. For a finite set of primes $P$ denote by $\gamma_{P}^{\mu}\in\mathcal{D}$ the element whose $p$-adic components are equal to those of $\mu$ for all $p\in P$ and equal to those of $\gamma$ for all other $p$. For example $\gamma_{\emptyset}^{\mu}=\gamma$ and $\gamma_{P}^{\mu}=\mu$ if $P$ contains all primes $p\mid N$. For $a\in\mathbb{Z}$ we have $a\gamma_{P}^{\mu}=(a\gamma)_{P}^{a\mu}$. We define
$v_{\gamma,\mu,P}\coloneqq\sum_{S\subset P}(-1)^{|S|}e^{\gamma_{S}^{\mu}}.$
Then we have
@@
Let $\mathcal{D}$ be a discriminant form of even signature, $P$ a finite set of prime numbers and $\gamma,\mu\in\mathcal{D}$. Assume that $\left(\prod_{p\in P}p\right)(\gamma-\mu)=0$, $\operatorname{q}(\gamma)=\operatorname{q}(\mu)\bmod 1$ and for all $p\in P$ assume $\gamma,\mu\not\in\mathcal{D}^{p}$ and if $2\in P$, also $\gamma,\mu\not\in\mathcal{D}^{2*}$. Let $s\in\mathbb{C}$ and $f\in\mathrm{S}_{k}(\mathcal{D})$ and assume that
$\prod_{p\in P}\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{rs}}\right)f$
converges. Then
$\langle\prod_{p\in P}\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{rs}}\right)f,v_{\gamma,\mu,P}\rangle=\langle f,v_{\gamma,\mu,P}\rangle.$
@@
We prove the statement by induction on $|P|$: Note that for $P=\emptyset$ there is nothing to prove.
So let $|P|\geq 1$ and assume that the assertion holds for all sets smaller than $P$. Let $p\in P$. We have
$v_{\gamma,\mu,P}=v_{\gamma,\mu_{\{p\}}^{\gamma},P\setminus\{p\}}-v_{\gamma_{\{p\}}^{\mu},\mu,P\setminus\{p\}}.$
The pairs $(\gamma,\mu_{\{p\}}^{\gamma})$ and $(\gamma_{\{p\}}^{\mu},\mu)$ satisfy the conditions of the corollary for the set of primes $P\setminus\{p\}$ and so by the induction hypothesis we have
$\langle\prod_{q\in P\setminus\{p\}}\left(\sum_{r=0}^{\infty}\frac{T(q^{2r})}{q^{rs}}\right)f,v_{\gamma,\mu,P}\rangle=\langle f,v_{\gamma,\mu,P}\rangle.$
In fact, all elements of the form $\gamma_{S}^{\mu}$ for $S\subset P$ satisfy the conditions of Proposition 3.4 with $\operatorname{q}(\gamma_{S}^{\mu})=\operatorname{q}(\gamma)=\operatorname{q}(\mu)\bmod 1$. Hence, we have for $r\geq 1$
$\langle T(p^{2r})f,v_{\gamma,\mu_{\{p\}}^{\gamma},P\setminus\{p\}}\rangle$ $=T^{\operatorname{q}(\gamma)}(p^{2r})\langle f,v_{p^{r}\gamma,p^{r}\mu_{\{p\}}^{\gamma},P\setminus\{p\}}\rangle\text{ and}$
$\langle T(p^{2r})f,v_{\gamma_{\{p\}}^{\mu},\mu,P\setminus\{p\}}\rangle$ $=T^{\operatorname{q}(\gamma)}(p^{2r})\langle f,v_{p^{r}\gamma_{\{p\}}^{\mu},p^{r}\mu,P\setminus\{p\}}\rangle.$
Furthermore, $p\gamma=p\gamma_{\{p\}}^{\mu}$ and $p\mu=p\mu_{\{p\}}^{\gamma}$, so that the right-hand sides are equal, thus
$\langle\left(\sum_{r=0}^{\infty}\frac{T(p^{2r})}{p^{rs}}\right)f,v_{\gamma,\mu,P}\rangle$ $=\langle f,v_{\gamma,\mu,P}\rangle+\langle\left(\sum_{r=1}^{\infty}\frac{T(p^{2r})}{p^{rs}}\right)f,v_{\gamma,\mu_{\{p\}}^{\gamma},P\setminus\{p\}}-v_{\gamma_{\{p\}}^{\mu},\mu,P\setminus\{p\}}\rangle$
$=\langle f,v_{\gamma,\mu,P}\rangle+\sum_{r=1}^{\infty}\frac{T^{\operatorname{q}(\gamma)}(p^{2r})}{p^{rs}}\langle f,v_{p^{r}\gamma,p^{r}\mu_{\{p\}}^{\gamma},P\setminus\{p\}}-v_{p^{r}\gamma_{\{p\}}^{\mu},p^{r}\mu,P\setminus\{p\}}\rangle$
$=\langle f,v_{\gamma,\mu,P}\rangle.$
@@
The case $x\in\mathbb{Z}$ is well-known (see e.g. [23, (7.1.9)]). For $x=\frac{p}{q}$ with $(p,q)=1$ write
@@
for elements $v,w,u\in\mathbb{C}[\mathcal{D}]$ and extend antilinearly in the second argument. The following proposition adapts [37, Proposition 1], which is originally due to Petersson, to the vector valued case.
@@
where $P_{\gamma,r}$ is the Poincaré series of index $(\gamma,r)$ as defined in [7]. There it is shown that